OpenStudy (anonymous):

A manufacturer finds that the revenue generated by selling x gadgets is given by the function R(x)=320x-.2x^2, where the revenue R(x) is measured in dollars. What is the max revenue?

OpenStudy (anonymous):

vertex formula

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have the possibilities if that would be any help

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

$ 800 $ 1600 $ 128000 $ 384000 $ 256000

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

128000 is the answer..

OpenStudy (anonymous):

okay.. i am right.. thanks for the practice.. 128000 i'll have to remember this for my physics test when determining the max height of a projectile

OpenStudy (anonymous):

1. recognize it is a parabola and use the vertex formula or take the function of its first derivative \[R(x)=320x-.2x^2 \] \[x = -b/2a \rightarrow -(320/(2(-.2)))\] x=800 plug in x to find R(x) \[R(800)=320x-.2x^2 =128000\]